the line segment joining the points 1,2 and -2,1 is divided by the line 3x+4y=7 in the ratio

Grade:11

1 Answers

Nishant Vora IIT Patna
7 years ago

Equation of line Joining (1,2) and (-2,1)

3y – 6 = x – 1
3y = x + 5
3x + 4y  = 7 (Solve and after that section formula)